# Quasi-Monte Carlo with one categorical variable

### v2

@@ 6 A toy example
    A plain RQMC estimator: $\hat{\mu}_{\mathrm{RQMC}}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})$ where $\bm{z}_{1},\dots,\bm{z}_{n}$ are the first $n$ points of a scrambled Sobol’ sequence in $[0,1]^{2}$.
@@
For each estimator we compute $R=500$ replicates at sample sizes $n=2^{m}$ for $3\leqslant m\leqslant 12$ and then graph the sample variances. We used nested uniform scrambling (Owen, 1995) to do the scrambling.
@@
Figure 2: Variance of the 5 estimators for the estimation of ${\mathbb{E}}(f(\bm{z}))$ versus the sample size $n$, for the toy integrand. Four of the curves are nearly parallel to reference lines of a given slope.
@@ 7 Saint-Venant flood model
We are motivated by engineering problems with substantial uncertainty. Great uncertainty often arises from weather variables or customer behavior. Our example is the Saint-Venant flood model from Limbourg and De Rocquigny (2010). A more precise flood model would be based on Navier-Stokes equations that have to be solved numerically. The Saint-Venant model is a simplification that we think is well suited for illustrative purposes.
@@
$Z_{v}$ Uniform ($49$,$51$) none
$Z_{m}$ Uniform ($54$,$56$) none
@@
This model is usually given with a symmetric triangular density for $D$ over the interval $[4990,5010]$. It is usually seen that uncertainty in $D$ makes very little difference. See for example Iooss and Lemaître (2015). Therefore we fix $D=5000$. Similarly, $B$ is commonly given a symmetric triangular distribution over $[295,305]$ and we fix it at $B=300$. We consider four distributions for the inputs $Q$, $K_{s}$, $Z_{v}$ and $Z_{m}$. With probability $0.95$ all four random variables are independently distributed with nominal distributions given in Table 2. With probability $0.02$, the distribution of $Q$ takes the adverse distribution in Table 2 while the other variables are nominal. With probability $0.02$, the distribution of $K_{s}$ is adverse and the others are nominal. With probability $0.01$ both $Q$ and $K_{s}$ are adverse.
The distribution of $K_{s}$ is customarily given as $\mathcal{N}(30,7.5^{2})$ and then truncated to a range such as $(0,\infty)$ in Limbourg and De Rocquigny (2010) or $(15,\infty)$ in Iooss (2011). We have chosen a nominal Gamma distribution, so we do not need any truncation. That Gamma distribution has mean $30$ but a lower variance than $7.5^{2}$. The adverse distribution for the Gamma distribution takes lower values which bring more flooding as $K_{s}$ appears in the denominator of $H$. The distribution of $Q$ is customarily given as Gumbel with mode $1013$ and scale $558$. This was found by a maximum likelihood fit, reported in Limbourg and De Rocquigny (2010). The Gumbel distribution has an extreme value justification, but it also gives negative values which are unphysical. We have replaced it by a Fréchet distribution which is always positive and is also an extreme value distribution. It has much heavier tails than the Gumbel which one might need to characterize extreme risks. Our nominal Fréchet distribution has a similar mean to the Gumbel distribution above. The adverse distribution triples the nominal one.
Figure 3: We plot variance versus sample size for $\hat{\mu}_{\mathrm{MC}}$, $\hat{\mu}_{\mathrm{RQMC}}$, $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$, $\hat{\mu}_{\mathrm{RQMC}(2)}$ and $\hat{\mu}_{\mathrm{RQMC}(L)}$ for the Saint-Venant flood depth (20). The sample size allocations for $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$, $\hat{\mu}_{\mathrm{RQMC}(2)}$, and $\hat{\mu}_{\mathrm{RQMC}(L)}$ used $\rho=2$.
@@
In Figure 4 we consider integer allocations with different values of $\rho$ using equation (14) to select $n_{\ell}$. For large $n$, the lowest variances are for $\rho=2$ and $3$ with curves that are nearly parallel with slope close to $-1.8$. The curve for $\rho=2$ which is also in Figure 3 is best at larger $n$. The case of $\rho=1$ has proportional allocation and may have an empty stratum for $n\leqslant 128$ (as $128*0.01<2$).
Figure 4: Variance of $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$ versus the sample size $n$ for $\rho\in\{1,2,3,+\infty\}$. The reference line has slope $-1.8$.
@@ 8 Discussion
We also tried a complex importance sampling problem from Hesterberg (1995) modeling reliability of oil inventory for an energy utility. The integrand was of the mixture importance sampling form (1). We found only a moderate improvement with RQMC giving roughly $O(n^{-1.15})$ variance. The example comes from Hesterberg (1995) and was also studied by Li et al. (2013). Here we only mention it to show that RQMC with mixture sampling will not always give an improvement that is practically meaningful. The relevant integrands were discontinuous and we speculate that they may have had important high order interactions.
@@
We have considered one categorical variable. Some problems may well have two or more categorical variables. Suppose that there are $K$ categorical variables where the $k$’th one has $L_{k}$ levels. Then we have the choice of whether to use $K$ different components of RQMC to make the allocations, or to use one component with $L=\prod_{k=1}^{K}L_{k}$ levels in it. Optimizing the choice between these two approaches is outside the scope of this article.
@@ Acknowledgments
We thank Neil Sloane for pointing out the connection to sequence A002572. Thanks to Tim Hesterberg and Wentao Li for help with the oil inventory problem. This work was supported by the National Science Foundation under grant DMS-2152780.
Appendix A Monotonocity of $I_{0}$
Here we prove Proposition 4.
@@ Appendix A Monotonocity of $I_{0}$
$h(x)$ $=\biggl{(}\,\sum_{\ell=1}^{L}\alpha_{\ell}^{2-2\rho/(\rho+1)-\rho x}\biggr{)}\biggl{(}\,\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)+x}\biggr{)}^{\rho}$
$=\biggl{(}\,\sum_{\ell=1}^{L}\alpha_{\ell}^{2/(\rho+1)-\rho x}\biggr{)}\biggl{(}\,\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)+x}\biggr{)}^{\rho}.$
@@
Their derivatives with respect to $x$ are $S_{+}(r+1)$ and $-\rho S_{-}(r+1)$, respectively. Also, $S_{+}(r,0)=S_{-}(r,0)$ which we write as $S(r)$.
Now we write
$h(x)=S_{-}(0)S_{+}(0)^{\rho}$
so
$h^{\prime}(x)=-\rho S_{-}(1)S_{+}(0)^{\rho}+\rho S_{-}(0)S_{+}(0)^{\rho-1}S_{+}(1).$
Then

### v3

@@ 6 A toy example
    A plain RQMC estimator: $\hat{\mu}_{\mathrm{RQMC}}=(1/n)\sum_{i=1}^{n}f(\bm{z}_{i})$ where $\bm{z}_{1},\dots,\bm{z}_{n}$ are the first $n$ points of a scrambled Sobol’ sequence in $0,1]^{2}$.
@@
For each estimator we compute $R=500$ replicates at sample sizes $n=2^{m}$ for $3\leqslant m\leqslant 12$ and then graph the sample variances. We used nested uniform scrambling (Owen 1995) to do the scrambling.
@@
Figure 2: Variance of the 5 estimators for the estimation of ${\mathbb{E}}(f(\bm{z}))$ versus the sample size $n$, for the toy integrand. The top to bottom ordering of the curves in the legend is the same as they have at $n=4096$. Four of the curves are nearly parallel to reference lines of a given slope.
@@ 7 Saint-Venant flood model
We are motivated by engineering problems with substantial uncertainty. Great uncertainty often arises from weather variables or customer behavior. Our example is the Saint-Venant flood model from Limbourg & De Rocquigny (2010). A more precise flood model would be based on Navier-Stokes equations that have to be solved numerically. The Saint-Venant model is a simplification that we think is well suited for illustrative purposes.
@@
$Z_{v}$ Uniform (4949,$51$) none
$Z_{m}$ Uniform (5454,$56$) none
@@
This model is usually given with a symmetric triangular density for $D$ over the interval $[4990,5010]$. It is usually seen that uncertainty in $D$ makes very little difference. See for example Iooss & Lemaître (2015). Therefore we fix $D=5000$. Similarly, $B$ is commonly given a symmetric triangular distribution over $[295,305]$ and we fix it at $B=300$. We consider four distributions for the inputs $Q$, $K_{s}$, $Z_{v}$ and $Z_{m}$. With probability $0.95$ all four random variables are independently distributed with nominal distributions given in Table 2. With probability $0.02$, the distribution of $Q$ takes the adverse distribution in Table 2 while the other variables are nominal. With probability $0.02$, the distribution of $K_{s}$ is adverse and the others are nominal. With probability $0.01$ both $Q$ and $K_{s}$ are adverse.
The distribution of $K_{s}$ is customarily given as $\mathcal{N}(30,7.5^{2})$ and then truncated to a range such as $(0,\infty)$ in Limbourg & De Rocquigny (2010) or $(15,\infty)$ in Iooss (2011). We have chosen a nominal Gamma distribution, so we do not need any truncation. That Gamma distribution has mean $30$ but a lower variance than $7.5^{2}$. The adverse distribution for the Gamma distribution takes lower values which bring more flooding as $K_{s}$ appears in the denominator of $H$. The distribution of $Q$ is customarily given as Gumbel with mode $1013$ and scale $558$. This was found by a maximum likelihood fit, reported in Limbourg & De Rocquigny (2010). The Gumbel distribution has an extreme value justification, but it also gives negative values which are unphysical. We have replaced it by a Fréchet distribution which is always positive and is also an extreme value distribution. It has much heavier tails than the Gumbel which one might need to characterize extreme risks. Our nominal Fréchet distribution has a similar mean to the Gumbel distribution above. The adverse distribution triples the nominal one.
Figure 3: We plot variance versus sample size for $\hat{\mu}_{\mathrm{MC}}$, $\hat{\mu}_{\mathrm{RQMC}}$, $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$, $\hat{\mu}_{\mathrm{RQMC}(2)}$ and $\hat{\mu}_{\mathrm{RQMC}(L)}$ for the Saint-Venant flood depth (15). The top to bottom ordering of the curves in the legend is the same as they have at $n=4096$. The sample size allocations for $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$, $\hat{\mu}_{\mathrm{RQMC}(2)}$, and $\hat{\mu}_{\mathrm{RQMC}(L)}$ used $\rho=2$.
@@
In Figure 4 we consider integer allocations with different values of $\rho$ using equation (9) to select $n_{\ell}$. For large $n$, the lowest variances are for $\rho=2$ and $3$ with curves that are nearly parallel with slope close to 1.8-1.8. The curve for $\rho=2$ which is also in Figure 3 is best at larger $n$. The case of $\rho=1$ has proportional allocation and may have an empty stratum for $n\leqslant 128$ (as 1280.01<2128*0.01<2).
Figure 4: Variance of $\hat{\mu}_{\mathrm{RQMC}(\mathrm{ADJ})}$ versus the sample size $n$ for $\rho\in\{1,2,3,+\infty\}$. The reference line has slope 1.8-1.8.
@@ 8 Discussion
for $\bm{x}_{i}\sim\sum_{\ell=1}^{L}\alpha_{\ell}p_{\ell}$.
We tried a complex importance sampling problem from Hesterberg (1995) modeling reliability of oil inventory for an energy utility. The estimator was of the mixture importance sampling form (16). An interesting aspect of mixture importance sampling is that it benefits from using $p_{\ell^{\prime}}(\bm{x}_{i})$ when $\bm{x}_{i}$ was drawn from $p_{\ell}$ for $\ell\neq\ell^{\prime}$. We found only a moderate improvement with RQMC giving roughly $O(n^{-1.15})$ variance. The example was also studied by Li et al. (2013). Here we only mention it to show that RQMC with mixture sampling will not always give an an extremely large improvement. The relevant integrands were discontinuous and we speculate that they may also have had important high order interactions.
@@
We have considered one categorical variable. Some problems will have two or more categorical variables. Suppose that there are $K$ categorical variables where the $k$’th one has $L_{k}$ levels. Then we have the choice of whether to use $K$ different components of RQMC to make the allocations, or to use one component with $L=\prod_{k=1}^{K}L_{k}$ levels in it. Optimizing the choice between these two approaches is outside the scope of this article.
@@ Acknowledgments
We thank Neil Sloane for pointing out the connection to sequence A002572. Thanks to Tim Hesterberg and Wentao Li for help with the oil inventory problem. We thank two anonymous referees and an associate editor for comments that helped us improve the paper.
[…]
This appendix contains the proofs of our propositions.
[…]
Proof.
Write $A=r/n+a$ and $B=s/n+b$ for integers $0\leqslant r\leqslant s<n$ and $a,b\in[0,1/n)$. Then $\beta=t/n+c$ for t=srt=s-r and 1/n<c<1/n-1/n<c<1/n. If $r=s$, then $0\leqslant n_{*}\leqslant 1$ and then (5) holds trivially.
For $s>r$ we may write
$A,B)=[A,(r+1)/n)\,\cup\,[(r+1)/n,s/n)\,\cup\,[s/n,s/n+b).$
Then $[A,B)$ contains at least sr1s-r-1 of the intervals $I_{k}=[(k-1)/n,k/n)$ for $k=1,\dots,n$ and it intersects with at most sr+1s-r+1 of them. Therefore $|n_{*}-(s-r)|\leqslant 1$. Now $\lceil n\beta\rceil=s-r+\lceil nc\rceil$ and so $s-r\geqslant\lceil n\beta\rceil-1$ establishing the lower limit in (5). Similarly $\lfloor n\beta\rfloor=s-r+\lfloor nc\rfloor$ so $s-r\leqslant\lfloor n\beta\rfloor+1$ establishing the upper limit in (5). ∎
[…]
Proof.
For $S_{\ell}$ defined at (3), sample $i\in S_{\ell}$ holds if and only if $B_{\ell-1}\leqslant z_{i1}<B_{\ell}$ or equivalently $\sum_{1\leqslant r<\ell}2^{-\kappa_{r}}\leqslant z_{i1}<\sum_{1\leqslant r\leqslant\ell}2^{-\kappa_{r}}$ which we can write as
Now consider a dyadic box in $0,1]^{s}$ of the form
$\prod_{j=1}^{s}\Bigl[\frac{c_{j}}{2^{k_{j}}},\frac{c_{j}+1}{2^{k_{j}}}\Bigr).$
[…]
Proof.
Its absolute value is at most
$\Bigl|\frac{n\beta_{\ell}}{n_{\ell}}-1\Bigr|\mathrm{Vol}([\bm{0},\bm{c}))+\frac{2n}{n_{\ell}}O(n^{-1}\log(n)^{s})$ $=O(n_{\ell}^{-1}\log(n_{\ell}/\beta_{\ell})^{s})$
$=O(n_{\ell}^{-1}\log(n_{\ell})^{s})$
[…]
@@
$h(x)$ $=\biggl(\,\sum_{\ell=1}^{L}\alpha_{\ell}^{2-2\rho/(\rho+1)-\rho x}\biggr)\biggl(\,\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)+x}\biggr)^{\rho}=\biggl(\,\sum_{\ell=1}^{L}\alpha_{\ell}^{2/(\rho+1)-\rho x}\biggr)\biggl(\,\sum_{k=1}^{L}\alpha_{k}^{2/(\rho+1)+x}\biggr)^{\rho}.$
@@
Their derivatives with respect to $x$ are $S_{+}(r+1)$ and $\rho S_{-}(r+1)$, respectively. Also, $S_{+}(r,0)=S_{-}(r,0)$ which we write as $S(r)$.
Now we write $h(x)=S_{-}(0)S_{+}(0)^{\rho}$ so $h^{\prime}(x)=-\rho S_{-}(1)S_{+}(0)^{\rho}+\rho S_{-}(0)S_{+}(0)^{\rho-1}S_{+}(1)$. Then
@@
[…]
Proof.
From equation (3.2)
[…]
Proof.
[…]
Proof.
The same argument applies to both criteria. We prove it for criterion 1. Fixing $\bm{n}$ and $\tilde{\bm{n}}$ we see that the maximum over $\bm{\tau}$ of $R_{1}(\bm{n}\!\mid\!\tilde{\bm{n}};\bm{\tau})$ arises when $\tau_{\ell}=0$ for all $\ell$ except $\ell^{*}(\bm{n},\tilde{\bm{n}})=\operatorname*{argmax}_{\ell}(\tilde{n}_{\ell}/n_{\ell})^{\rho/2}$ with $\tau_{\ell^{*}(\bm{n},\tilde{\bm{n}})}$ then taking any value $\lambda>0$.
Now, we can go back to the initial problem and rewrite (13) as:
after noting that the factor $\alpha_{\ell^{*}(\bm{n},\tilde{\bm{n}})}\tau_{\ell^{*}(\bm{n},\tilde{\bm{n}})}^{1/2}$ cancels from numerator and denominator. To maximize this ratio over $\tilde{\bm{n}}$ one takes $\tilde{\bm{n}}$ to be the vector with value NL+1N-L+1 in component $\ell^{*}=\operatorname*{argmin}_{\ell}n_{\ell}$ and ones elsewhere. Then the minimax problem becomes $\min_{\bm{n}\in\Delta_{N}}\bigl((N-L+1)/\min_{1\leqslant\ell\leqslant N}n_{\ell}\bigr)^{\rho/2}$. The solution is to maximize the minimum value of $n_{\ell}$. We do that by making $\min_{\ell}n_{\ell}=\lfloor N/L\rfloor$. ∎
[…]
Our model for sample allocation is that we start with $n_{\ell}=1$ for $\ell=1,\dots,L$. Then we repeatedly increment $n_{\ell^{*}}$ by one for $\ell^{*}=\arg\max_{\ell}\alpha_{\ell}^{2}(n_{\ell}^{-\rho_{\ell}}-(n_{\ell}+1)^{-\rho_{\ell}})$ because this gives the stratum that most reduces the criterion. For our RQMC estimate to be consistent, we require $\min_{\ell}n_{\ell}\to\infty$, and so we assume that all of the $n_{\ell}$ are large enough to allow a Taylor approximation below. We also ignore ties in this analysis.
Using $(n_{\ell}+1)^{-\rho_{\ell}}=n_{\ell}^{-\rho_{\ell}}-\rho_{\ell}n_{\ell}^{-(\rho_{\ell}+1)}+O(n_{\ell}^{-\rho_{\ell}-2})$, we simplify the selection rule to $\ell^{*}=\arg\max_{\ell}\alpha_{\ell}^{2}\rho_{\ell}/n_{\ell}^{\rho_{\ell}+1}.$ Then for $1\leqslant\rho_{\ell}<\rho_{\ell^{\prime}}\leqslant 3$ under our greedy allocation $n_{\ell^{\prime}}$ will only get a new point if
$n_{\ell^{\prime}}<\Bigl(\frac{\alpha_{\ell^{\prime}}^{2}\rho_{\ell^{\prime}}}{\alpha_{\ell}^{2}\rho_{\ell}}n_{\ell}^{\rho_{\ell}+1}\Bigr)^{1/(\rho_{\ell^{\prime}}+1)}$
and so we expect $n_{\ell^{\prime}}\lesssim n_{\ell}^{(\rho_{\ell}+1)/(\rho_{\ell^{\prime}}+1)}$, where $\lesssim$ denotes inequality up to a multiplicative constant. For the most extreme case, with $\rho_{\ell}=1$ and $\rho_{\ell^{\prime}}=3$, we get $n_{\ell^{\prime}}\lesssim n_{\ell}^{1/2}$.